# Motivic Homotopy Groups of Spheres and Free Summands of Stably Free Modules

### v2

@@ 1.2. Remarks and conventions
Some of our later results, notably Proposition 4.3 and Theorem 4.4, are concerned with maps that are isomorphisms modulo uniquely divisible subgroups of the source. We devote some attention to the homological algebra of such morphisms in Appendix A.
@@
Some of our sources, particularly [8], [10], [14] and other works on the motivic Adams spectral sequence, work over the ground field $\mathbb{C}$. Using [21], for the purposes of calculating $\pi_{s,w}(\mathbb{1}^{\wedge}_{p})$, we can safely replace $\mathbb{C}$ by an arbitrary algebraically closed field $k$ of characteristic 0. We ultimately wish to apply our calculations, in Section 5, in the case where $k=\bar{\mathbb{Q}}$, which is the most general available to us. It imposes essentially no extra cost to work throughout over a general algebraically closed field $k$ of characteristic 0.
The calculations of homotopy groups in this paper all hold over an algebraically closed field $k$ of characteristic $\ell>0$, at the possible cost of replacing $\pi_{s,w}(X)$ by $\mathbb{Z}[\ell^{-1}]\otimes_{\mathbb{Z}}\pi_{s,w}(X)$. In fact, the calculations here, and knowledge of the Milnor–Witt $K$-theory, suffice to determine the analogous calculations over even more general fields. We refer the interested reader to [5, Theorem 1.1].
@@ 2. Ext-Completions
We make no claims about the originality of this section, the results of which are elementary and, in many cases, implicit in [24]. It is included here for reference and to set notation.
@@ 2.1. Definitions
The notation above concerns the group as a whole. Given an element $a\in A$, we say that $a$ is $I$-torsion if it is $n$-torsion for some integer $n$ that is a product of primes in $I$; it is $n$-divisible (in $A$) if the equation $nx=a$ can be solved for $x$ in $A$, and uniquely $n$-divisible the equation has a unique solution; $a$ is (uniquely) $I$-divisible if it is (uniquely) $n$-divisible for all integers $n$ that are products of primes in $I$.
@@
Suppose $F:\mathrm{Ab}^{\text{op}}\to\mathrm{Ab}$ is a contravariant additive functor. From the previous definitions, it is immediate that
  • If $A$ is uniquely $I$-divisible, then $F(A)$ is uniquely $I$-divisible;
  • If $F$ is left exact (e.g., $F(-)=\operatorname{Hom}(-,B)$ for some $B$), and if $A$ is $I$-torsion-free, then $F(A)$ is $I$-torsion-free.
Write $I^{\prime}=\mathcal{P}\setminus I$. We write $\mathbb{Z}[I^{-1}]$ to denote the subring of $\mathbb{Q}$ consisting of numbers that may be written as $a/b$ where $a$ is an integer and $b$ is a product of primes in $I$. Note that $\mathbb{Z}[\emptyset^{-1}]=\mathbb{Z}$ and $\mathbb{Z}[\mathcal{P}^{-1}]=\mathbb{Q}$. We define $\mathbb{Z}/\left(I^{\infty}\right)$ as the quotient group:
$0\to\mathbb{Z}\to\mathbb{Z}[I^{-1}]\to\mathbb{Z}/\left(I^{\infty}\right)\to 0.$
The group $\mathbb{Z}/\left(I^{\infty}\right)$ is torsion, and the torsion order of each element a product of primes in $I$. In the case where $I=\{p\}$ is a singleton, we generally write $\mathbb{Z}[p^{-1}]$ and $\mathbb{Z}/\left(p^{\infty}\right)$. If $I=\mathcal{P}$ is the set of all prime numbers, then we will write $\mathbb{Q}/\mathbb{Z}$ for $\mathbb{Z}/[\mathcal{P}^{\infty}]$.
About these groups, we observe
  • $\mathbb{Z}[I^{-1}]$ is uniquely $I$-divisible.
  • $\mathbb{Z}/\left(I^{\infty}\right)$ is $I$-divisible, and uniquely $I^{\prime}$-divisible: for $I$, this follows from the same fact for $\mathbb{Z}[I^{-1}]$. Suppose $q$ is a prime not in $I$. For any $x\in\mathbb{Z}/\left(I^{\infty}\right)$, we can find some product $n$ of primes in $I$ such that $nx=0$. Then $q$ divides $n^{q-1}-1$, so that $\frac{1-n^{q-1}}{q}$ is an integer. If we multiply $\frac{1-n^{q-1}}{q}x$ by $q$, we obtain $x$, so that $\times q$ is indeed surjective. Injectivity of $\times q$ follows from this: for any $y\in\mathbb{Z}[I^{-1}]$, the product $qy$ is an integer if and only if $y$ is.
We deduce that, for any abelian group $A$:
@@
Definition 2.1 ([6]).
@@
We have observed above that the $\operatorname{Ext}$-completion consists of nontorsion or elements whose torsion order is a product of primes in $I$. From the short exact sequence
(2) $0\to\mathbb{Z}\to\mathbb{Z}[I^{-1}]\to\mathbb{Z}/\left(I^{\infty}\right)\to 0$
we deduce that there is a natural transformation (a connecting homomorphism)
@@ 2.2. Surjectivity of completion
Let $A$ be an abelian group and $I$ a set of prime numbers. Suppose $\operatorname{Ext}(\mathbb{Z}/\left(I^{\infty}\right),A)$ is a torsion group. Then the completion $A\to\operatorname{Ext}(\mathbb{Z}/\left(I^{\infty}\right),A)$ is surjective.
@@
It suffices to show that the map $j$ is 0. The target is uniquely $I$-divisible, so that $\times n^{-1}$ is a well defined automorphism of $\operatorname{Ext}(\mathbb{Z}[I^{-1}],A)$ whenever $n$ is a product of primes in $I$. By hypothesis the source consists of torsion elements, and orders of these elements are products of primes in $I$. For any such $a\in\operatorname{Ext}(\mathbb{Z}/\left(I^{\infty}\right),A)$, we can find some $n$ for which $na=0$. Then $j(na)=0$, but where $nj(a)=0$ implies $j(a)=0$. ∎
@@
Consider $\operatorname{Ext}(C,A)$, on which $\times n$ acts as an isomorphism, since $A$ is uniquely $I$-divisible, and as 0, since $C$ is $n$-torsion. It must be the case that $\operatorname{Ext}(C,A)=0$. The Yoneda interpretation of $\operatorname{Ext}$ implies the sequence splits. ∎
@@
and a splitting map $s$. In this diagram, $j$ is the completion map, the image of $g$ is exactly the subgroup of $I$-divisible elements in $A$, and the image of $s$ is the $I$-torsion subgroup of $A$.
@@ 3. The sphere spectrum over $k$
We continue to work over an algebraically closed field $k$ of characteristic 0. The previous results on completions apply well to the motivic sphere spectrum, $\mathbb{1}$. If $p$ is a prime number and $X$ is a motivic spectrum, then [21, §3] constructs the $p$-completion $X\to X^{\wedge}_{p}$. There are exact sequences
(6) $0\to\operatorname{Ext}(\mathbb{Z}/\left(p^{\infty}\right),\pi_{r,s}(X))\overset{i}{\to}\pi_{r,s}(X^{\wedge}_{p})\overset{j}{\to}\operatorname{Hom}(\mathbb{Z}/\left(p^{\infty}\right),\pi_{r-1,s}(X))\to 0.$
Relying on [21], we know that the groups $\pi_{r,s}(X^{\wedge}_{p})$ are independent of the particular choice of $k$. These groups are intensively studied, and the following result is well known to the experts in the field. We learned the proof from Dan Isaksen and Zhouli Xu.
@@
Let $s$ be an integer different from 0 and let $w$ be an integer. As $p$ ranges over the prime numbers, the groups $\pi_{s,w}(\mathbb{1}^{\wedge}_{p})$ are finite, and almost all of them are 0.
@@
For any given weight $w$, there is an Adams–Novikov spectral sequence converging to $\pi_{*,w}(\mathbb{1}^{\wedge}_{p})$: this is constructed for $p=2$ in [12] and in [25] for odd primes (see also [26]). The $E_{2}$-page of this spectral sequence is isomorphic to a bigraded subgroup of the $E_{2}$-page of the classical Adams–Novikov spectral sequence. This follows from [12, Equation (36)] in the case of $p=2$, and from [25, Prop. 2.4.4(2)]. The classical vanishing results [19, Cor. 2.1, Cor 3.1] imply that for each stem $s$, there are only finitely many primes $p$ for which these groups are not 0, and for each such prime $p$, there are only finitely many associated nonzero groups on the $E_{2}$-page and these are themselves finite abelian groups. Since the $E_{\infty}$-page of the motivic Adams–Novikov spectral sequence is a subquotient of the $E_{2}$-page, it follows that $\pi_{s,w}(\mathbb{1}^{\wedge}_{p})$ has a finite filtration whose associated graded groups are finite. In particular, $\pi_{s,w}(\mathbb{1}^{\wedge}_{p})$ is a finite abelian group. ∎
Theorem 3.2.
Let $s,w$ be integers satisfying $s\neq 0$. Then the comparison map
is split surjective, and the kernel is the subgroup of divisible elements. Additionally, if $s\neq-1$, then the kernel is isomorphic to the motivic cohomology group $\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Z}(-w))$.
@@
We remark that the motivic cohomology groups $\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Z}(-w))$ are uniquely divisible except when $w\leq 0$ and s=0,1s=0,-1. This is a consequence of the long exact sequence
and the norm-residue isomorphism theorem ([11, Thm. 1.8]) for the algebraically closed field $k$. As a result, rationalization induces an isomorphism
$\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Z}(-w))\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Q}(-w)),\text{\quad provided $s\neq 0,-1$.}$
We use this to replace the rational cohomology by the integral at the end of the proof.
@@
For a fixed $p$, in the exact sequence (6), we remark that $\operatorname{Hom}(\mathbb{Z}/\left(p^{\infty}\right),\pi_{s-1,w}(\mathbb{1}))$ is a torsion-free quotient of a torsion group, so it vanishes, and therefore $\operatorname{Ext}(\mathbb{Z}/\left(p^{\infty}\right),\pi_{s,w}(\mathbb{1}))\cong\pi_{s,w}(\mathbb{1}^{\wedge}_{p})$. From this, Corollary 2.3 tells us that the comparison
@@
Since this group is finite, Proposition 2.4 applies, saying that the completion map is surjective. When $s\neq-1$, we have also seen that $\operatorname{Hom}(\mathbb{Z}/\left(p^{\infty}\right),\pi_{s+1,w}(\mathbb{1}))=0$, so that Proposition 2.6 applies, telling us the completion is split surjective and the kernel is the subgroup of divisible elements.
In the case of s=1s=-1, a different argument is required. In this case, the Adams and Adams–Novikov spectral sequences imply (see [10] and [25, Prop. 2.4.8]) that $\pi_{-1,w}(\mathbb{1}^{\wedge}_{p})=0$ for all primes $p$. Therefore the completion map is split surjective for trivial reasons. We also see that $\operatorname{Ext}(\mathbb{Z}/\left(p^{\infty}\right),\pi_{-1,w}(\mathbb{1}))=0$, by short exact sequence (6). We may, however, have a nontrivial presentation arising from exact sequence (4):
@@
For the rest of this proof, we assume $s\neq-1$. The comparison map belongs in a (split) short exact sequence:
$0\to K_{s,w}\to\pi_{s,w}(\mathbb{1})\to\prod_{p\in\mathcal{P}}\pi_{s,w}(\mathbb{1}^{\wedge}_{p})\to 0$
where the kernel $K_{s,w}$ is uniquely divisible, i.e., a $\mathbb{Q}$-vector space, and the cokernel is finite. Therefore, applying $\mathbb{Q}\otimes_{\mathbb{Z}}$- to the short exact sequence yields an isomorphism
(7) $K_{s,w}\cong\mathbb{Q}\otimes_{\mathbb{Z}}\pi_{s,w}(\mathbb{1}).$
Except in the case where $s=w$, an isomorphism $\mathbb{Q}\otimes_{\mathbb{Z}}\pi_{s,w}(\mathbb{1})\cong\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Q}(-w))$ is contained in [2, Thm. 3.7].
We finally consider the case of $s=w$. Here we know the groups $\pi_{s,s}(\mathbb{1})\cong{{\mathbf{K}}}^{\mathsf{MW}}_{-s}(k)$, from [18, Thm. 1.23, Cor. 1.25]. There is a presentation of ${{\mathbf{K}}}^{\mathsf{MW}}_{-s}(k)$ as a product of the Milnor $K$-theory group ${{\mathbf{K}}}^{\mathsf{M}}_{-s}(k)$ with a power $I^{-s}(k)$ of the fundamental ideal of the Witt ring of the field, fibred over $\mathbf{k}^{\mathsf{M}}_{-s}(k)$ the mod-$2$ reduction of the Milnor $K$-theory: see [18, Lem. 3.10]. Over an algebraically closed field, the fundamental ideal vanishes, as does the mod-$2$ Milnor $K$-theory. We are left with an isomorphism ${{\mathbf{K}}}^{\mathsf{MW}}_{-s}(k)\cong{{\mathbf{K}}}^{\mathsf{M}}_{-s}(k)$. The isomorphism $\mathbb{Q}\otimes_{\mathbb{Z}}{{\mathbf{K}}}^{\mathsf{M}}_{-s}(k)\cong\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Q}(-s))$ follows from [17, Ch. 5]. ∎
@@
First, suppose $s\neq-1$ and $w\geq-1$. Then Theorem 3.2 tells us the map is surjective and the kernel is a motivic cohomology group $\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Z}(-w))$. This group vanishes by [17, Ch. 4].
Now suppose s=1s=-1 and $w\geq 0$. Here $\pi_{-1,w}(\mathbb{1})=\pi_{-1-w+w\alpha}(\mathbb{1})=0$ by Morel’s calculations [18, Cor. 6.43], so that surjectivity of the comparison map implies isomorphism. ∎
@@
$\pi_{s,w}(\mathbb{1})\to\pi_{s}(\mathbb{S})$
and similarly for the $p$-completions.

### v3

@@ 1.1. Outline
In our later results, notably in Proposition 4.3 and Theorem 1.5, we are concerned with maps that are isomorphisms modulo uniquely divisible subgroups of the source. We consider the homological algebra of such morphisms in Appendix A.
1.2. The ground field
@@
Some of our sources, particularly [undefg], [undefj], [undefq] and other works on the motivic Adams and Adams–Novikov spectral sequences, work over the ground field $\mathbb{C}$. For the purposes of calculating $\pi_{s,w}(\mathbb{1}^{\wedge}_{p})$, we can use [undefad] to replace $\mathbb{C}$ by an arbitrary algebraically closed field $k$ of characteristic 0. We ultimately wish to apply our calculations, in Section 5, in the case where $k=\bar{\mathbb{Q}}$, which is the most general available to us. It imposes essentially no extra cost to work throughout over a general algebraically closed field $k$ of characteristic 0.
@@ 1.3. Acknowledgements
We thank an anonymous referee for a thoughtful and constructive reading of the paper.
@@ 2. Ext-Completions
We make no claims about the originality of this section, the results of which are elementary and, in many cases, implicit in [undefag]. It is included here for reference and to set notation.
@@ 2.1. Definitions
The notation above concerns the group as a whole. Given an element $a\in A$, we say that $a$ is $I$-torsion if it is $n$-torsion for some integer $n$ that is a product of primes in $I$; it is $n$-divisible (in $A$) if the equation $nx=a$ can be solved for $x$ in $A$, and uniquely $n$-divisible if the equation has a unique solution; $a$ is (uniquely) $I$-divisible if it is (uniquely) $n$-divisible for all integers $n$ that are products of primes in $I$.
@@
Suppose $F:\mathrm{Ab}^{\text{op}}\to\mathrm{Ab}$ is a contravariant additive functor. From the previous definitions, it is immediate that if $A$ is uniquely $I$-divisible, then $F(A)$ is uniquely $I$-divisible.
Write
$I^{\prime}=\mathcal{P}\setminus I.$
We use $\mathbb{Z}[I^{-1}]$ to denote the subring of $\mathbb{Q}$ consisting of numbers that may be written as $a/b$ where $a$ is an integer and $b$ is a product of primes in $I$. Note that $\mathbb{Z}[\emptyset^{-1}]=\mathbb{Z}$ and $\mathbb{Z}[\mathcal{P}^{-1}]=\mathbb{Q}$. We define $\mathbb{Z}/\left(I^{\infty}\right)$ as the quotient group in the short exact sequence:
(3) $0\to\mathbb{Z}\to\mathbb{Z}[I^{-1}]\to\mathbb{Z}/\left(I^{\infty}\right)\to 0.$
The group $\mathbb{Z}/\left(I^{\infty}\right)$ is torsion, and the torsion order of each element is a product of primes in $I$. In the case where $I=\{p\}$ is a singleton, we generally write $\mathbb{Z}[p^{-1}]$ and $\mathbb{Z}/\left(p^{\infty}\right)$. If $I=\mathcal{P}$ is the set of all prime numbers, then we will write $\mathbb{Q}/\mathbb{Z}$ for $\mathbb{Z}/[\mathcal{P}^{\infty}]$.
Lemma 2.1.
The group $\mathbb{Z}[I^{-1}]$ is uniquely $I$-divisible. The group $\mathbb{Z}/\left(I^{\infty}\right)$ is $I$-divisible, and uniquely $I^{\prime}$-divisible
Proof.
The first claim, unique $I$-divisibility of $\mathbb{Z}[I^{-1}]$, is elementary.
The $I$-divisibility of $\mathbb{Z}/\left(I^{\infty}\right)$ follows from the same fact for $\mathbb{Z}[I^{-1}]$. Our last claim, about $I^{\prime}$-divisibility of $\mathbb{Z}/\left(I^{\infty}\right)$, is vacuously true if $I=\mathcal{P}$ and is trivially true if $I^{\prime}=\mathcal{P}$. We assume that $\emptyset\neq I\neq\mathcal{P}$.
Suppose $q$ is a prime not in $I$. For any $x\in\mathbb{Z}/\left(I^{\infty}\right)$, we can find some product $n$ of primes in $I$ such that $nx=0$. Then $q$ divides $n^{q-1}-1$, so that $\frac{1-n^{q-1}}{q}$ is an integer. If we multiply $\frac{1-n^{q-1}}{q}x$ by $q$, we obtain $x$, so that $\times q$ is indeed surjective. Injectivity of $\times q$ follows from this: for any $y\in\mathbb{Z}[I^{-1}]$, the product $qy$ is an integer if and only if $y$ is. ∎
Since $\operatorname{Hom}$, $\operatorname{Ext}$ are additive functors, we immediately deduce the following.
Corollary 2.2.
For any abelian group $A$:
  1. (1)
@@
Proof.
Points 1 and 2 hold since $\operatorname{Hom}(-,A)$ and $\operatorname{Ext}(-,A)$ are additive and the groups $\mathbb{Z}[I^{-1}]$ and $\mathbb{Z}/\left(I^{\infty}\right)$ are uniquely $I$- and $I^{\prime}$-divisible respectively.
Since uniquely $I^{\prime}$-divisible implies $I^{\prime}$-torsion free, to establish 3 we need only show that $\operatorname{Hom}(\mathbb{Z}/\left(I^{\infty}\right),A)$ is $I$-torsion-free. The left-exact functor $\operatorname{Hom}(-,A)$ converts the epimorphism $\mathbb{Z}[I^{-1}]\to\mathbb{Z}/\left(I^{\infty}\right)$ to an injection, so that $\operatorname{Hom}(\mathbb{Z}/\left(I^{\infty}\right),A)$ is isomorphic to a subgroup of an $I$-torsion-free group, establishing the result. ∎
Definition 2.3 ([undefd]).
@@
We have observed above that the $\operatorname{Ext}$-completion consists of nontorsion or elements whose torsion order is a product of primes in $I$. From the short exact sequence (3) we deduce that there is a natural homomorphism
@@ 2.2. Surjectivity of completion
Let $A$ be an abelian group and $I$ a set of prime numbers. Suppose $\operatorname{Ext}(\mathbb{Z}/\left(I^{\infty}\right),A)$ is a torsion group. Then the $I$-completion $A\to\operatorname{Ext}(\mathbb{Z}/\left(I^{\infty}\right),A)$ is surjective.
@@
It suffices to show that the map $j$ is 0. The target is uniquely $I$-divisible, so that $\times n^{-1}$ is a well defined automorphism of $\operatorname{Ext}(\mathbb{Z}[I^{-1}],A)$ whenever $n$ is a product of primes in $I$. By hypothesis the source consists of torsion elements, and the orders of these elements are products of primes in $I$. For any $a\in\operatorname{Ext}(\mathbb{Z}/\left(I^{\infty}\right),A)$, we can find some $n$ for which $na=0$. Then $0=nj(a)$ by linearity, whereupon $j(a)=0$. ∎
@@
Since $A$ is uniquely $I$-divisible, multiplication by $n$ is an automorphism of $\operatorname{Ext}(C,A)$. On the other hand, since $C$ is $n$-torsion, multiplication by $n$ on $\operatorname{Ext}(C,A)$ is the 0-map. We conclude that $\operatorname{Ext}(C,A)=0$. The Yoneda interpretation of $\operatorname{Ext}$ implies the sequence splits. ∎
@@
and a splitting map $s$. In this diagram, $j$ is the $I$-completion map, the image of $g$ is exactly the subgroup of $I$-divisible elements in $A$, and the image of $s$ is the $I$-torsion subgroup of $A$.
@@ 3. The sphere spectrum over $k$
We continue to work over an algebraically closed field $k$ of characteristic 0. The previous results on completions apply well to the motivic sphere spectrum, $\mathbb{1}$. If $p$ is a prime number and $X$ is a motivic spectrum, then [undefad, §3] constructs the $p$-completion $X\to X^{\wedge}_{p}$, and assures us that there are exact sequences
(7) $0\to\operatorname{Ext}\Big(\mathbb{Z}/\left(p^{\infty}\right),\pi_{r,s}(X)\Big)\overset{i}{\to}\pi_{r,s}(X^{\wedge}_{p})\overset{j}{\to}\operatorname{Hom}(\mathbb{Z}/\left(p^{\infty}\right),\pi_{r-1,s}(X))\to 0.$
Relying on [undefad], we know that the groups $\pi_{r,s}(X^{\wedge}_{p})$ are independent of the particular choice of $k$. These groups are intensively studied, and the following result is well known to the experts in the field. We learned the outline of the proof from Dan Isaksen and Zhouli Xu.
@@
Let $s$ be an integer different from 0 and let $w$ be an integer. Let $p$ be a prime number. The group $\pi_{s,w}(\mathbb{1}^{\wedge}_{p})$ is a finite abelian $p$-group, and it vanishes if s<2p3s<2p-3.
@@
For any given weight $w$, there is an Adams–Novikov spectral sequence converging to $\pi_{*,w}(\mathbb{1}^{\wedge}_{p})$, constructed in [undefo] for $p=2$ and in [undefy, Thm. 3.1] for odd primes. It is a convergent spectral sequence of trigraded groups
$\mathrm{E}_{2}^{n,t,w}=\operatorname{Ext}^{n,t,w}_{\mathsf{BP}_{*,*}\mathsf{BP}}(\pi_{*,*}\mathsf{BP},\pi_{*,*}\mathsf{BP})\Longrightarrow\pi_{t-n,w}(\mathbb{1}^{\wedge}_{p}),$
where $\mathsf{BP}$ is the $p$-completion of the $p$-local motivic Brown–Peterson spectrum, which was first constructed in [undefaj] and [undefn]. The differentials satisfy $d_{r}:\mathrm{E}_{r}^{n,t,w}\to\mathrm{E}_{r}^{n+r,t+r-1,w}$, and in particular, the weight-grading $w$ is preserved by the differentials so that we may view this as a family of spectral sequences, one for each $w\in\mathbb{Z}$. We fix $w$ for the rest of this proof.
Comparison of the motivic $\mathrm{E}_{2}$-page with the $\mathrm{E}_{2}$-page of the classical Adams–Novikov spectral sequence is given by [undefo, Equation (36)] in the case of $p=2$, and by [undefah, Prop. 2.4.4(2)] when $p$ is odd. The comparison is given by complex realization, and on the level of groups yields
(8) $\operatorname{Ext}^{n,t,w}_{\mathsf{BP}_{*,*}\mathsf{BP}}(\pi_{*,*}\mathsf{BP},\pi_{*,*}\mathsf{BP})\cong\begin{cases}\operatorname{Ext}^{n,t}_{\mathsf{BP}^{\mathrm{top}}_{*}\mathsf{BP}^{\mathrm{top}}}(\pi_{*}\mathsf{BP}^{\mathrm{top}},\pi_{*}\mathsf{BP}^{\mathrm{top}})&\text{if $w\leq t/2$;}\\ 0&\text{otherwise.}\end{cases}$
Here the notation $\mathsf{BP}^{\mathrm{top}}$ is used for the $p$-complete classical Brown–Peterson spectrum. We will write $\operatorname{Ext}^{n,t,w}$ for these groups for the sake of brevity.
Similarly, write $\operatorname{Ext}^{n,t}$ for $\operatorname{Ext}^{n,t}_{\mathsf{BP}^{\mathrm{top}}_{*}\mathsf{BP}^{\mathrm{top}}}(\pi_{*}\mathsf{BP}^{\mathrm{top}},\pi_{*}\mathsf{BP}^{\mathrm{top}})$. A well-known algebraic presentation of these groups, which may be found in [undefaa, Prop. 3.16], shows that these groups are finitely generated modules over the ring of $p$-adic integers $\mathbb{Z}^{\wedge}_{p}$. This presentation also makes it clear that the groups vanish unless $n\geq 0$ and $t\geq 0$. The statement [undefab, Thm. 5.2.1(a)] implies that they are abelian $p$-groups, and therefore finite, except when $n=t=0$. Sparseness results [undefx, Cor. 3.1] imply that these groups also vanish when $t\not\equiv 0\pmod{2p-2}$ and when t<2n(p1)t<2n(p-1). Beware that [undefx] uses $Q_{p}$” to denote the ring of $p$-adic integers, and argues using $\mathsf{MU}$ rather than $\mathsf{BP}$, but the implication for $\operatorname{Ext}^{n,t}$ is immediate. The results of this paragraph also apply to the subgroups (up to isomorphism) $\operatorname{Ext}^{n,t,w}$ of $\operatorname{Ext}^{n,t}$.
The vanishing results imply two things: first, there are only finitely many nonzero groups $\operatorname{Ext}^{n,t,w}$ for any given value of $s$, owing to the constraint $0\leq 2(p-1)n\leq t$; second, the smallest nonzero value of tnt-n for which there is a group $\operatorname{Ext}^{n,t}$ that is not 0 is tn=2p3t-n=2p-3, attained by the pair (n,t)=(1,2p2)(n,t)=(1,2p-2). This is reflected in the well-known fact that $p$-torsion arises in the classical stable homotopy groups for the first time at $\pi_{2p-3}(\mathbb{S})$.
The $E_{\infty}$-page of the motivic Adams–Novikov spectral sequence converging to $\pi_{t-n,w}(\mathbb{1}^{\wedge}_{p})$ is a subquotient of the $E_{2}$-page, and it follows that when $t-n\neq 0$, the group $\pi_{t-n,w}(\mathbb{1}^{\wedge}_{p})$ has a finite filtration whose associated graded groups are finite abelian $p$-groups. In particular, $\pi_{t-n,w}(\mathbb{1}^{\wedge}_{p})$ is a finite abelian group when $t-n\neq 0$. Additionally, this group vanishes when tn<2p3t-n<2p-3, except when tn=0t-n=0. ∎
Let us write $\mathsf{M}\mathbb{Z}$ for the motivic cohomology spectrum. There exists a unit map $\mathbb{1}\to\mathsf{M}\mathbb{Z}$, arising as a unit map of an adjunction between the stable homotopy category and a triangulated category of motives as in [undefe, §5.3.35]. This induces a unit homomorphism $\eta:\pi_{s,w}(\mathbb{1})\to\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Z}(-w))$.
We are now in a position to prove our first main theorem, which we restate for the convenience of the reader. See 1.1
@@
For a fixed $p$, in the exact sequence (7), the group $\operatorname{Hom}(\mathbb{Z}/\left(p^{\infty}\right),\pi_{s-1,w}(\mathbb{1}))$ is a torsion-free quotient of a torsion group. It vanishes, yielding an isomorphism $\operatorname{Ext}(\mathbb{Z}/\left(p^{\infty}\right),\pi_{s,w}(\mathbb{1}))\cong\pi_{s,w}(\mathbb{1}^{\wedge}_{p})$. From this, Corollary 2.5 tells us that the comparison
@@
Since this group is finite, Proposition 2.6 applies, saying that the completion map
$\pi_{s,w}(\mathbb{1})\to\operatorname{Ext}(\mathbb{Q}/\mathbb{Z},\pi_{s,w}(\mathbb{1}))$
is surjective. When $s\neq-1$, we have also seen that $\operatorname{Hom}(\mathbb{Z}/\left(p^{\infty}\right),\pi_{s+1,w}(\mathbb{1}))=0$, so that Proposition 2.8 applies, telling us the completion is split surjective and the kernel is the subgroup of divisible elements. Furthermore, the kernel is isomorphic to $\operatorname{Hom}(\mathbb{Q},\pi_{s,w}(\mathbb{1}))$, so is a $\mathbb{Q}$-vector space in its own right.
In the case of s=1s=-1, a different argument is required. In this case, Proposition 3.1 tells us that that $\pi_{-1,w}(\mathbb{1}^{\wedge}_{p})=0$ for all primes $p$. Therefore the completion map is split surjective for trivial reasons. We also see that $\operatorname{Ext}(\mathbb{Z}/\left(p^{\infty}\right),\pi_{-1,w}(\mathbb{1}))=0$, by short exact sequence (7). We may, however, have a nontrivial presentation arising from exact sequence (5):
@@
For the rest of this proof, we assume $s\neq-1$. There is a diagram
[…]
where the row is an exact sequence, so that $i:K_{s,w}\to\pi_{s,w}(\mathbb{1})$ denotes inclusion of the kernel of the completion map. It remains to prove that $\eta\circ i$ is an isomorphism.
The group $K_{s,w}$ was observed to be a $\mathbb{Q}$-vector space, and the cokernel of $i$ is finite, so that applying $\mathbb{Q}\otimes_{\mathbb{Z}}$- to (9) yields a diagram
[…]
in which the row is exact.
A proof that
$\mathrm{id}_{\mathbb{Q}}\otimes_{\mathbb{Z}}\eta\colon\mathbb{Q}\otimes_{\mathbb{Z}}\pi_{s,w}(\mathbb{1})\to\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Q}(-w))$
is an isomorphism is outlined in [undefy, Thm. 6.2], following work of Morel (unpublished) and [undefe, §§5.3, 16.2]. Strictly speaking, these references assert only that there is an abstract isomorphism, but from [undefe, § 5.3.35] one sees that this abstract isomorphism is induced by the unit map. In particular $\mathrm{id}_{\mathbb{Q}}\otimes(\eta\circ i)$ is an isomorphism.
We claim that the motivic cohomology groups $\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Z}(-w))$ are uniquely divisible, i.e., admit the structure of $\mathbb{Q}$-vector spaces, except when $w\leq 0$ and s=0,1s=0,-1. This is a consequence of the long exact sequence
and the fact that
(10) $\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Z}/(n)(-w))\cong\begin{cases}\mathbb{Z}/(n)&\text{ if $s=0$ and $w\leq 0$;}\\ 0&\text{ otherwise}.\end{cases}$
The isomorphisms in (10) are justified by:
  • The vanishing of motivic cohomology for nonpositive weights, i.e., when $w\geq 0$ [undefu, Cor. 4.2];
  • Dimensional vanishing, i.e., when $s<w$ [undefu, Thm. 3.5];
  • The Beilinson–Lichtenbaum version of the norm-residue isomorphism, [undefk, Thm. 1.8(b)], which says that
[…]
This proves our claim.
Therefore the map $\eta\circ i:K_{s,w}\to\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Z}(-w))$ is a homomorphism between $\mathbb{Q}$ vector spaces that yields an isomorphism after application of $\mathbb{Q}\otimes_{\mathbb{Z}}$-. In particular, it is itself an isomorphism. ∎
@@
First, suppose $s\neq-1$ and $w\geq-1$. Then Theorem 1.1 tells us the map is surjective and the kernel is the motivic cohomology group $\mathrm{H}^{-s}(\operatorname{Spec}k;\mathbb{Z}(-w))$. This group vanishes by [undefu, Ch. 4].
Now suppose s=1s=-1 and $w\geq 0$. Here $\pi_{-1,w}(\mathbb{1})=\pi_{-1-w+w\alpha}(\mathbb{1})=0$ by Morel’s calculations [undefv, Cor. 6.43], so that surjectivity of the comparison map implies isomorphism. ∎
@@
$\pi_{s,w}(\mathbb{1})\to\pi_{s}(\mathbb{S}),$
by virtue of the realization functors [undefz, Thm. A.45]. Recall from [undefad] that the $p$-completion $X^{\wedge}_{p}$ is constructed by means of Bousfield localization with respect to the motivic Moore spectrum. One construction of this spectrum is $\mathbb{1}/(p)=\mathbb{S}/(p)\wedge\mathbb{1}$, where $\mathbb{S}/(p)$ denotes the classical Moore spectrum. Complex realization of $\mathbb{1}/(p)$ yields $\mathbb{S}/(p)$, since this is a monoidal functor the realization of $\mathbb{1}$ is $\mathbb{S}$, the monoidal unit. The theory of Bousfield localization, e.g., [undefl, Thm. 3.3.20], implies that the two functors, $p$-completion and complex realization, commute up to equivalence. There are commutative squares
$\pi_{s,w}(\mathbb{1})$$\pi_{s}(\mathbb{S})$$\pi_{s,w}(\mathbb{1}^{\wedge}_{p})$$\pi_{s}(\mathbb{S}^{\wedge}_{p}).$}